# SimPO: Simple Preference Optimization with a Reference-Free Reward

### v2

@@ 5 Related Work
RLHF is a technique that aligns large language models with human preferences and values [16, 97, 60, 7]. The classical RLHF pipeline typically comprises three phases: supervised fine-tuning [96, 74, 31, 19, 46, 23, 79, 14, 82], reward model training [30, 58, 15, 54, 35, 48], and policy optimization [68, 4]. Proximal Policy Optimization (PPO) [68] is a widely used algorithm in the third stage of RLHF. The RLHF framework is also widely applied to various applications, such as mitigating toxicity [47, 3], ensuring safety [22], enhancing helpfulness [75], serching and navigating the web [59], and improving model reasoning abilities [34]. Recently, [13] has highlighted challenges across the whole RLHF pipeline from preference data collection to model training. Further research has also demonstrated that RLHF can lead to biased outcomes, such as verbose outputs from the model [26, 69, 81].
@@
Given that online preference optimization algorithms are complex and difficult to optimize [95, 67], researchers have been exploring more efficient and simpler alternative offline algorithms. Direct Preference Optimization (DPO) [64] is a notable example. However, the absence of an explicit reward model in DPO limits its ability to sample preference pairs from the optimal policy. To address this, researchers have explored augmenting preference data using a trained SFT policy [92] or a refined SFT policy with rejection sampling [57], enabling the policy to learn from data generated by the optimal policy. Further studies have extended this approach to an iterative training setup, continuously updating the reference model for optimization with either the most recent policy model or by generating preference pairs scored and selected by the evolving policy model [44, 65, 88, 25, 83]. In this work, we focus exclusively on offline settings, avoiding any iterative training processes.
@@
A variety of preference optimization objectives have been proposed besides DPO. Ranking objectives allow for comparisons among more than two instances [24, 56, 70, 87]. Another line of work explores simpler preference optimization objectives that do not rely on a reference model [40, 85], similar to SimPO. [8] proposes a method to jointly optimize instructions and responses, finding it effectively improves DPO. [93] focuses on post-training extrapolation between the SFT and the aligned model to further enhance model performance. In this work, we compare SimPO to a series of offline algorithms, including RRHF [87], SLiC-HF [92], DPO [64], IPO [6], CPO [84], KTO [27], ORPO [40], and R-DPO [62], and find that SimPO can outperform these methods in both efficiency and performance. Recently, [73] proposed a generalized preference optimization framework that unifies different offline algorithms, and SimPO can be seen as a special case of this framework.
6 Discussion
Conclusion.
In this work, we propose SimPO, a simple and effective preference optimization algorithm that consistently outperforms existing approaches across various training setups. By aligning the reward function with the generation likelihood and introducing a target reward margin, SimPO eliminates the need for a reference model and achieves strong performance without exploiting the length bias. Extensive analysis demonstrates that the key designs in SimPO are crucial and validates the efficiency and effectiveness of SimPO.
Limitations and future work.
First, despite the empirical success and intuitive motivation of SimPO, we lack a theoretical and rigorous understanding of why it works. Moreover, the introduction of a target reward margin requires us to tune an additional hyperparameter, future work could explore how to determine the optimal margin automatically. Second, SimPO is an offline preference algorithm and does not leverage iterative training or other orthogonal techniques. Future work could explore the integration of SimPO with these methods to further enhance model performance. Third, our experiments focus solely on the evaluation of helpfulness, neglecting other crucial aspects of model behavior such as safety, honesty, and fairness. Investigating the generalization of SimPO to these behaviors in future studies is important. Finally, we observe a performance drop on some downstream tasks, especially on the math benchmarks. We provide a more detailed discussion of the limitations and future directions in the next section.
Limitations
@@ Conclusion.
Acknowledgments
Appendix A Implementation Details
@@
CPO [84] $-\log\sigma\left(\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{l}|x)\right)-\lambda\log\pi_{\theta}(y_{w}|x)$ $\alpha=1.0,\,\,\beta\in[0.01,0.05,0.1]$
@@
R-DPO [62] $-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\left(\alpha|y_{w}|-\alpha|y_{l}|\right)\right)$ $\alpha\in[0.05,0.1,0.5,1.0]$
@@
Llama3-Base 2.0 1.0 6e-7
Llama3-Instruct 2.5 1.4 1e-6
@@
For AlpacaEval 2, we use a sampling decoding strategy to generate responses, with a temperature of 0.7 for the Mistral-Base setting following zephyr-7b-beta, [^12] https://github.com/tatsu-lab/alpaca_eval/blob/main/src/alpaca_eval/models_configs/zephyr-7b-beta/configs.yaml a temperature of 0.5 for the Mistral-Instruct setting following Snorkel-Mistral-PairRM-DPO, and a temperature of 0.9 for both Llama3 settings. [^13] We grid search the temperature hyperparameter for the Llama3-Base setting with DPO over ${0.1,0.3,0.5,0.7,0.9}$, and fix it for all different methods. For Arena-Hard, we use the default greedy decoding for all settings and methods. For MT-Bench, we follow the official decoding configuration which defines different sampling temperatures for different categories.
@@
Appendix B Downstream Task Evaluation
@@
Appendix C Standard Deviation of AlpacaEval 2 and Arena-Hard
@@
Appendix D Generation Length Analysis
@@
We observe that SimPO has a shorter generation length compared to DPO in the Llalam3-Instruct case but exhibits a higher generation length in other settings, with up to 26% longer responses on AlpacaEval 2. Conversely, SimPO only increases length by only around 5% on Arena-Hard compared to DPO. It is fair to say that the generation length heavily depends on the evaluation benchmark. A stronger indicator is that SimPO consistently achieves a higher length-controlled win rate on AlpacaEval 2 compared to the raw win rate, demonstrating minimal exploitation of response length.
@@
Appendix E Gradient Analysis
@@
Appendix F Qualitative Analysis
@@
In Figure 11, we present an example where Llama3-Instruct generates a more detailed and well-formatted answer compared to the baseline model, and as well as the Llama3-Base-SimPO model. Given the question: What language does Argentina people speak? Llama3-Base-SimPO only gives a very brief answer. GPT-4-Preview-1106 gives a more detailed answer in explaining how the Argentina Spanish differs from standard Spanish. However, the answer is not well formatted and a bit hard to parse. Llama3-Instruct-SimPO gives a detailed and well-formatted answer, which is easier to read and understand, and offers sufficient details.
@@ 3. Follow-up actions:
Figure 11: A case study on AlpacaEval 2 demonstrates that Llama3-Instruct, trained with SimPO, provides a better formatted and more detailed answer than both Llama3-Base, also trained with SimPO, and the baseline model GPT-4-1106-Preview. This illustrates how the instruction setting typically outperforms the base setting.
@@
Llama3-Base SimPO $\to$ The official language spoken in Argentina is Spanish, specifically Argentine Spanish, which shares similarities with other varieties of Spanish spoken in neighboring countries like Uruguay, Chile, and Paraguay.
Llama3-Instruct SimPO $\to$ In Argentina, the official language is Spanish. Specifically, the dialect spoken in Argentina is known as Rioplatense Spanish, which is a variation of Spanish with some unique characteristics and influences from Italian, German, and other immigrant languages.
@@
Appendix G Llama3-Instruct v0.2 (Jul 7, 2024)
Table 12: Results of Llama3-Instruct (8B) setting, utilizing preference labels annotated by a stronger reward model (ArmoRM [80], we term it as version 0.2).
Method Llama3-Instruct (8B)
@@
In this section, we update the Llama3-Instruct setting, primarily by utilizing a stronger reward model to annotate our generated preference dataset.
@@
We use the following hyperparameters for SimPO under the Llama3 Instruct v0.2 setting: $\beta=10$ and $\gamma=3$. The other hyperparameters (e.g., learning rate, batch size, max sequence lengths) are kept the same as the original Llama3-8B-Instruct setting.
@@
With a strong SFT model like Llama3-Instruct-8B, and as the preference optimization data quality improves, the differences between algorithms become less pronounced. For instance, DPO achieved a similar win rate as SimPO in terms of raw win rate, and DPO, IPO, and R-DPO all exhibited comparable raw win rates on Arena-Hard. However, SimPO maintains an advantage by producing shorter sequences, resulting in a significantly better LC win rate on AlpacaEval 2.
@@
$\mathcal{L}_{\text{SimPO w/ SFT}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)-\lambda\log\pi_{\theta}(y_{w}|x)\right].$
@@
Appendix H Applying Length Normalization and Target Reward Margin to DPO (Jul 7, 2024)
Since the release of the paper, we have had inquiries from researchers about whether the key design elements of SimPO – length normalization and target reward margin – could benefit DPO. By doing so, we will derive the following two objectives:

### v3

@@ 5 Related Work
RLHF is a technique that aligns large language models with human preferences and values [18, 102, 62, 7]. The classical RLHF pipeline typically comprises three phases: supervised fine-tuning [101, 76, 33, 21, 48, 25, 82, 15, 86], reward model training [32, 60, 16, 56, 37, 50], and policy optimization [70, 4]. Proximal Policy Optimization (PPO) [70] is a widely used algorithm in the third stage of RLHF. The RLHF framework is also widely applied to various applications, such as mitigating toxicity [3, 49, 97], ensuring safety [24], enhancing helpfulness [78, 83], searching and navigating the web [61], and improving model reasoning abilities [36]. Recently, [13] has highlighted challenges across the whole RLHF pipeline from preference data collection to model training. Further research has also demonstrated that RLHF can lead to biased outcomes, such as verbose outputs from the model [28, 71, 85].
@@
Given that online preference optimization algorithms are complex and difficult to optimize [100, 69], researchers have been exploring more efficient and simpler alternative offline algorithms. Direct Preference Optimization (DPO) [66] is a notable example. However, the absence of an explicit reward model in DPO limits its ability to sample preference pairs from the optimal policy. To address this, researchers have explored augmenting preference data using a trained SFT policy [96] or a refined SFT policy with rejection sampling [59], enabling the policy to learn from data generated by the optimal policy. Further studies have extended this approach to an iterative training setup, by continuously updating the reference model with the most recent policy model or generating new preference pairs at each iteration [27, 46, 67, 87, 92]. In this work, we focus exclusively on offline settings, avoiding any iterative training processes.
@@
A variety of preference optimization objectives have been proposed besides DPO. Ranking objectives allow for comparisons among more than two instances [26, 58, 72, 91]. Another line of work explores simpler preference optimization objectives that do not rely on a reference model [42, 89], similar to SimPO. [8] proposes a method to jointly optimize instructions and responses, finding it effectively improves DPO. [98] focuses on post-training extrapolation between the SFT and the aligned model to further enhance model performance. In this work, we compare SimPO to a series of offline algorithms, including RRHF [91], SLiC-HF [96], DPO [66], IPO [6], CPO [88], KTO [29], ORPO [42], and R-DPO [64], and find that SimPO can outperform them in both efficiency and performance. Recently, [75] proposed a generalized preference optimization framework unifying different offline algorithms, and SimPO can be seen as a special case.
6 Conclusion
In this work, we propose SimPO, a simple and effective preference optimization algorithm that consistently outperforms existing approaches across various training setups. By aligning the reward function with the generation likelihood and introducing a target reward margin, SimPO eliminates the need for a reference model and achieves strong performance without exploiting the length bias. Extensive analysis demonstrates that the key designs in SimPO are crucial and validates the efficiency and effectiveness of SimPO. A detailed discussion of the limitations can be found in Appendix A.
Acknowledgments
Appendix A Limitations
@@
Appendix B Implementation Details
@@
CPO [88] $-\log\sigma\left(\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{l}|x)\right)-\lambda\log\pi_{\theta}(y_{w}|x)$ $\lambda=1.0,\,\,\beta\in[0.01,0.05,0.1]$
@@
R-DPO [64] $-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}+\left(\alpha|y_{w}|-\alpha|y_{l}|\right)\right)$ $\alpha\in[0.05,0.1,0.5,1.0]$
@@
Llama-3-Base 2.0 1.0 6e-7
Llama-3-Instruct 2.5 1.4 1e-6
@@
For AlpacaEval 2, we use a sampling decoding strategy to generate responses, with a temperature of 0.7 for the Mistral-Base setting following zephyr-7b-beta, [^13] https://github.com/tatsu-lab/alpaca_eval/blob/main/src/alpaca_eval/models_configs/zephyr-7b-beta/configs.yaml a temperature of 0.5 for the Mistral-Instruct setting following Snorkel-Mistral-PairRM-DPO, and a temperature of 0.9 for both Llama 3 settings. [^14] We grid search the temperature hyperparameter for the Llama-3-Base setting with DPO over ${0.1,0.3,0.5,0.7,0.9}$, and fix it for all different methods. For Arena-Hard, we use the default greedy decoding for all settings and methods. For MT-Bench, we follow the official decoding configuration which defines different sampling temperatures for different categories.
@@
Appendix C Downstream Task Evaluation
@@
Appendix D Standard Deviation of AlpacaEval 2 and Arena-Hard
@@
Appendix E Generation Length Analysis
@@
We observe that SimPO has a shorter generation length compared to DPO in the Llama-3-Instruct case but exhibits a higher generation length in other settings, with up to 26% longer responses on AlpacaEval 2. Conversely, SimPO only increases length by only around 5% on Arena-Hard compared to DPO. It is fair to say that the generation length heavily depends on the evaluation benchmark. A stronger indicator is that SimPO consistently achieves a higher length-controlled win rate on AlpacaEval 2 compared to the raw win rate, demonstrating minimal exploitation of response length.
@@
Appendix F Gradient Analysis
@@
Appendix G Qualitative Analysis
@@
In Figure 11, we present an example where Llama-3-Instruct generates a more detailed and well-formatted answer compared to the baseline model, and as well as the Llama-3-Base-SimPO model. Given the question: What language does Argentina people speak? Llama-3-Base-SimPO only gives a very brief answer. GPT-4-Preview-1106 gives a more detailed answer in explaining how the Argentina Spanish differs from standard Spanish. However, the answer is not well formatted and a bit hard to parse. Llama-3-Instruct-SimPO gives a detailed and well-formatted answer, which is easier to read and understand, and offers sufficient details.
@@ 3. Follow-up actions:
Figure 11: A case study on AlpacaEval 2 demonstrates that Llama-3-Instruct, trained with SimPO, provides a better formatted and more detailed answer than both Llama-3-Base, also trained with SimPO, and the baseline model GPT-4-1106-Preview. This illustrates how the instruction setting typically outperforms the base setting.
@@
Llama-3-Base SimPO $\to$ The official language spoken in Argentina is Spanish, specifically Argentine Spanish, which shares similarities with other varieties of Spanish spoken in neighboring countries like Uruguay, Chile, and Paraguay.
Llama-3-Instruct SimPO $\to$ In Argentina, the official language is Spanish. Specifically, the dialect spoken in Argentina is known as Rioplatense Spanish, which is a variation of Spanish with some unique characteristics and influences from Italian, German, and other immigrant languages.
@@
Appendix H Llama-3-Instruct v0.2 (Jul 7, 2024)
Table 12: Results of Llama-3-Instruct (8B) setting, utilizing preference labels annotated by a stronger reward model (ArmoRM [84], we term it as version 0.2).
Method Llama-3-Instruct (8B)
@@
In this section, we update the Llama-3-Instruct setting, primarily by utilizing a stronger reward model to annotate our generated preference dataset.
@@
We use the following hyperparameters for SimPO under the Llama-3 Instruct v0.2 setting: $\beta=10$ and $\gamma=3$. The other hyperparameters (e.g., learning rate, batch size, max sequence lengths) are kept the same as the original Llama-3-8B-Instruct setting.
@@
With a strong SFT model like Llama-3-8B-Instruct, and as the preference optimization data quality improves, the differences between algorithms become less pronounced. For instance, DPO achieved a similar win rate as SimPO in terms of raw win rate, and DPO, IPO, and R-DPO all exhibited comparable raw win rates on Arena-Hard. However, SimPO maintains an advantage by producing shorter sequences, resulting in a significantly better LC win rate on AlpacaEval 2.
@@
$\mathcal{L}_{\text{SimPO w/ SFT}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)+\lambda\log\pi_{\theta}(y_{w}|x)\right].$
@@
Appendix I Applying Length Normalization and Target Reward Margin to DPO (Jul 7, 2024)
Since the release of the paper, we have had inquiries from researchers about whether the key design elements of SimPO—length normalization and target reward margin—could benefit DPO. By doing so, we will derive the following two objectives:
@@ Appendix I Applying Length Normalization and Target Reward Margin to DPO (Jul 7, 2024)
Appendix J Applying SimPO to Gemma 2 Models (Sept 16, 2024)
Performance degradation on other benchmarks for Llama-3-SimPO checkpoints.
After releasing the Llama-3-SimPO checkpoints, we received extensive feedback about performance degradation on benchmarks measuring specific capabilities, such as MMLU and GSM8K. To investigate this issue, we continued training the Llama-3-8B-Instruct model with different learning rates, as reported in Table 16. We find that using a higher learning rate results in a stronger model in chat-oriented benchmarks, at the cost of catastrophic forgetting on GSM8K and MMLU. [^16] We evaluate the zero-shot performance of the models on GSM8K and MMLU using the ZeroEval repository which adopts a unified setup. With a smaller learning rate, the model’s performance on chat benchmarks is slightly worse, but its performance on GSM8K and MMLU is better retained. This demonstrates a trade-off between chat-oriented benchmarks and other benchmarks when continuing training from a strong instruction-tuned model.
Table 16: Results on AlpacaEval 2, ZeroEval GSM, and ZeroEval MMLU when continuing training from Llama-3-8B-Instruct with different learning rates. * indicates the released checkpoint.
Model AlpacaEval 2 ZeroEval ZeroEval
LC (%) GSM (%) MMLU (%)
Llama-3-8B-Instruct 26.0 78.5 61.7
SimPO (lr=4e-7) 38.8 77.9 62.6
SimPO (lr=5e-7) 44.6 77.0 62.3
SimPO (lr=1e-6)* 53.7 57.4 54.9
Applying SimPO to Gemma 2 models presents a different trend.
We evaluate SimPO using Google’s recently released Gemma-2-9B-it model [77], which represents a strong open-source model. For training data, we generate up to 5 responses per prompt from the UltraFeedback dataset [23] and use the ArmoRM model [84] to annotate preferences between responses. We compare our SimPO against a DPO-trained variant, both fine-tuned from the Gemma-2-9B-it base model. As shown in Table 17, SimPO demonstrates superior performance on chat benchmarks like AlpacaEval 2 and Arena-Hard while maintaining the model’s original zero-shot capabilities on tasks like GSM8K and MMLU. Notably, we find that varying the learning rate during fine-tuning has minimal impact on the model’s performance. These results suggest an underlying property difference between the Llama-3 checkpoints and the Gemma 2 checkpoints, and might be worth further investigation.
Gemma-2-9B-it-SimPO significantly improved the ranking of the Gemma-2-9B-it model on Chatbot Arena.
During the development stage, we relied solely on automated metrics to evaluate the model’s performance. To determine if these metrics aligned with real user preferences, we submitted our best-performing model, Gemma-2-9B-it-SimPO, to the Chatbot Arena leaderboard hosted by LMSYS [17]. We find that our model improved the original Gemma-2-9B-it ranking from 36th to 25th, making the SimPO variant the top-ranked $<$10B model on the Chatbot Arena leaderboard based on real user votes as of September 16th, 2024.
Table 17: Benchmark performance of Gemma-2-9B trained with DPO and SimPO on UltraFeedback (responses regenerated with Gemma-2-9B-it, following the same dataset construction process as Llama-3-Instruct (8B) described in Section 3). SimPO results in better instruction following performance than DPO without degrading math abilities (GSM) or general knowledge (MMLU) of the original model. * indicates the released checkpoint.
Model Instruction Following Capabilities
AlpacaEval 2 LC Arena-Hard ZeroEval ZeroEval
(%) (%) GSM (%) MMLU (%)
Gemma-2-9B-it 51.1 40.8 87.4 72.7
Gemma-2-9B-DPO 67.8 58.9 88.5 72.2
Gemma-2-9B-SimPO (lr=6e-7) 71.7 58.3 88.3 72.2
Gemma-2-9B-SimPO (lr=8e-7)* 72.4 59.1 88.0 72.2
Gemma-2-9B-SimPO (lr=1e-6) 71.0 58.3 87.4 71.5